OpenStudy (anonymous):

Help please? Solve by the elimination method. 4a + 4b =8 -4a + b =37 Is there a solution (type an ordered pair), infinitely many solutions, or no solutions?

OpenStudy (anonymous):

4(a+b) = 8 a+b = 4 subtract in second equation: -5a = 33 a = 33/-5 Therefore b = 8.6

OpenStudy (anonymous):

4a + 4b =8 + -4a + b =37 ------------------ 5b = 45 --- --- 5 5 ------------------ b = 9

OpenStudy (anonymous):

how do I type an ordered pair for these equations?

OpenStudy (anonymous):

There must be a few solutions

OpenStudy (anonymous):

Well that one I got incorrect, so I need to do a similar excercise, what i'm looking for is to be shown how to do these step by step so I can learn the process myself, can you help me? Here is the next equation.. 4a + 5b =13 -4a +b =17

OpenStudy (anonymous):

b=9 4a + 4b =8 4a+4(9)=8 4a=-28 a=-7 Ordered pair is (-7,9)

OpenStudy (anonymous):

Can you help me rework the next equation that I listed? 4a + 5b =13 -4a +b =17

OpenStudy (anonymous):

4a + 5b =13 + -4a +b =17 ----------------- 6b=30 b=5 4a+5(5)=13 4a+25=13 4a=-12 a=-3 (-3,5)

OpenStudy (anonymous):

Thank you Grazes & Violy. I appreciate the help.

OpenStudy (anonymous):

np.